Please help me solve this equation: x + 27/12 = 1/4x (the division symbol means it is a fraction)
Answer by stanbon(75887)

27/12 = 1/4x
-----------
Cross-Multiply to get:
(4x)(27) = 12
---
Divide by 4:
27x = 3
x = 3/27
x = 1/9
============
Cheers,
Stan H.